RC Meeting: August 25th, 2020

From clas12-run
Jump to: navigation, search

Back to Run Group F | Daily Run Group F Coordination Meetings

12:00PM RC Meeting (https://bluejeans.com/7576835804)

RC:


2020

March
19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30,31
July
31
August
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21 22 23 24, 25, 26, 27, 28, 29, 30, 31
September
1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20 21

Summary of Aug. 25th

Owl shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3833671
  2. Issues: Increasingly frequent RTPC trips. Even increasing the current limit on GEM2 to 8.5 µA doesn't help much.
  3. Recorded data: 26.4 M events on D2

Day shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3833932
  2. Issues: 3 hours of downtime for various accelerator improvements. We ran several tests to see if we could reduce the backgrounds: Lambertson sweep, moving the collimator, moving beam at 2C21 while keeping it fixed at 2H01. Nothing helped. Still seeing RTPC trips. Rebooting HV crate and swapping cables didn't help with the current in GEM2. Set GEM2 trip current to 10 µA. Beam rather unstable most of shift.
  3. Recorded data: RTPC pedestal run, runs with different currents 20nA - 100 nA. 3.8 M events total.

Swing shift

  1. Shift summary by the shift expert: https://logbooks.jlab.org/entry/3834284
  2. Issues: Slowly inching our way back up to 200 nA.
  3. Recorded data: 50 M

Whole day summary

  1. Total recorded data today: 80 M events
  2. Part-II total accumulated data on D2: 966 M
  3. Concerns: Backgrounds on beam counters and FD remain, but RTPC seems to be back to normal.


Tomorrow daily RC meeting will be at usual time 12:00 PM using the link https://bluejeans.com/7576835804/

Recorded data numbers

Number of events recorded